You have $150,000 to invest today. You found a deal that will pay you $300,000 if you wait 20 years before you withdraw the money? What is the implied rate of interest in you invest and wait 20 years?

Group of answer choices

3.2%

3.5%

5.1%

6.2%

6.8%

Please help urgently.
Answer #1

Ans 3.5%

FV = Future Value
PV = Present Value
r = rate of interest
n= no of period
FV/ PV = (1 + r )^n
300000/150000 = (1 + r%)^20
2 = (1 + r%)^20
r = 3.5%